Circular No. 7819

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ar IN NGC 3746 W. D. Li, University of California at Berkeley, reports the LOTOSS (cf.

IAUC 7514) discovery of an apparent supernova (mag ∼ 16.5) on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Feb. 3.5 and 4.5 UT. SN 2002ar is located at α = 11^{h}37^{m}43^{s}.92, δ = +22^{o}00′34′′.4 (equinox 2000.0), which is 3′′.3 east and 0′′.5 south of the nucleus of NGC 3746. A KAIT image taken on Jan. 18.5 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002as IN UGC 3418 Li also reports the LOTOSS discovery of an apparent supernova (mag

∼ 17.8) found by D. Weisz on an unfiltered KAIT image taken on Feb. 4.3 UT. SN 2002as is located at α = 6^{h}12^{m}25^{s}.05, δ = +44^{o}26′16′′.8 (equinox 2000.0), which is 11′′.3 west and 3′′.2 south of the nucleus of UGC 3418. The new object is confirmed at about the same brightness in an earlier KAIT image taken on Jan. 30.4. A KAIT image taken on Jan. 21.4 showed nothing at this position (limiting mag ∼ 19.5).

SUPERNOVA 2002ao IN UGC 9299 A. Gal-Yam and O. Shemmer, Tel Aviv University, write: “A full

reduction of a spectrum (range 400–780 nm) of SN 2002ao, obtained with the Wise Observatory 1-m telescope (+ FOSC spectrograph) on Jan. 31 by J. Dann, confirms the findings reported on IAUC 7810 from an earlier spectrum. The spectrum is quite blue, with the continuum rising by at least a factor of 2 between 700 and 500 nm. Strong absorption troughs are observed at 570 and 630 nm, and an emission feature around 590 nm is also seen. The blue continuum and strong absorption at 570 nm are not typical of early spectra of type-Ic supernovae (e.g., Matheson et al. 2001, A.J. 121, 1648), and they are not apparent in the spectrum of the type-Ic SN 1999bz referred to on IAUC 7815. The resemblance of SN 2002ao spectra to earlytime counterparts of the type-IIb SN 1996cb persists, making this type a likely identification for SN 2002ao. We encourage further spectroscopic observations of this object at later epochs.”

CORRIGENDA On IAUC 7773, ‘Supernova 2001ib’, line 16, for 20001ib read 2001ib On IAUC 7794, ‘Novae in M31’, line 17, for supernovae read novae

2002 February 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green